Find the accumulated amount A if the principal P = RM11,000 is invested at the interest rate of r = 5% per year for t = 5.5 years, compounded quarterly.
Answer by mananth(16946)

Find Amount
Principal P = 11000
Amount= A
years=t 5.50
compounded 4 times a year n
Rate = 5.00 0.05 0.0125
Amount = P*((n+r)/n)^n*t

Amount = = 11000 *( 1 + 0.0125 )^ 4 * 5.50
Amount = 11000 *( 1 + 0.0125 )^ 22
11000 *( 1.0125 )^ 22.00
Amount = 14457.17